Since he was 23 years old, Ben has been depositing $200 at the end of each month into a tax-free retirement account earning interest at the rate of 5.5%/year compounded monthly. Larry, who is the same age as Ben, decided to open a tax-free retirement account 5 years after Ben opened his. If Larry's account earns interest at the same rate as Ben's, determine how much Larry should deposit each month into his account so that both men will have the same amount of money in their accounts at age 65. (Round your answer to the nearest cent.) $
Answer #1

Calculating Future Value for Ben,

using TVM Calculation,

FV = [PV = 0, PMT = 200, N = 504, I = 0.055/12]

FV = $393,660.90

Calculating Monthly Deposit for Larry,

Using TVM calculation,

PMT = [PV = 0, FV = 393,660.90, I = 0.055/12, N = 444]

PMT = $272.68